A.A scientist measures the standard enthalpy change for the following reaction to be -2932.6 kJ :

2C2H6(g) + 7 O2(g)ebab9d75-6e9a-4c80-9bcd-b491b668f097.gif4CO2(g) + 6 H2O(g)

Based on this value and the standard enthalpies of formation for the other substances, the standard enthalpy of formation of CO2(g) is  kJ/mol.

B.A scientist measures the standard enthalpy change for the following reaction to be -138.9 kJ :

H2(g) + C2H4(g)ebab9d75-6e9a-4c80-9bcd-b491b668f097.gifC2H6(g)

Based on this value and the standard enthalpies of formation for the other substances, the standard enthalpy of formation of C2H4(g) is  kJ/mol.
Answer #1

Consider reaction, 2 C2H6 (g) + 7 O 2 (g) \rightarrow 4 CO 2 (g) + 6 H2O (g)  

The standard enthalpy change of reaction is given as

phpjTV1Rp.png r H 0 = phpjyELxE.pngphpXLdR9d.png H 0 f ( products ) -  phpe4m2PI.pngphp1NOs4N.png H 0 f ( reactants )

php5I6etK.pngphpSDfXlw.png H 0 f (products) = 4 phpNmo5NR.png H 0 f CO 2(g) +6 phpvVOk3l.png H 0 f H2O (g)

php5I6etK.pngphpSDfXlw.png H 0 f (products) = 4 phpNmo5NR.png H 0 f CO 2(g) + 6 ( -241.82 kJ /mol) =  4 phpNmo5NR.png H 0 f CO 2(g) - 1450.092 kJ / mol

phpaR7SqB.png H 0f (reactants) =2 phpxFwLMO.png H 0 f C2H6 (g) + 7 php2qKz93.png H 0 f O 2(g)

phpaR7SqB.png H 0f (reactants) = 2 (-84.0 kJ/ mol ) + 7 ( 0 kJ /mol) = - 168 kJ / mol

From above calculated values , we can write

- 2932.6 kJ / mol = (  4 phpNmo5NR.png H 0 f CO 2(g) - 1450.092 kJ / mol ) - ( - 168 kJ / mol )

- 2932.6 kJ / mol = (  4 phpNmo5NR.png H 0 f CO 2(g) - 1450.092 kJ / mol ) + 168 kJ / mol  

(  4 phpNmo5NR.png H 0 f CO 2(g) - 1450.092 kJ / mol ) = - 2932.6 kJ / mol - 168 kJ / mol = - 3100.6 kJ / mol

(  4 phpNmo5NR.png H 0 f CO 2(g) - 1450.092 kJ / mol ) = - 3100.6 kJ / mol

4 phpNmo5NR.png H 0 f CO 2(g) = - 3100.6 kJ / mol + 1450.092 kJ / mol = - 1650.51 kJ /mol

phpNmo5NR.png H 0 f CO 2(g) = - 1650.51 kJ /mol / 4 = - 412.6 kJ / mol

ANSWER : Standard enthalpy of formation of CO 2 = - 412.6 kJ / mol

PART 2

Consider reaction, C2H4 (g) + H 2 (g) \rightarrow C 2H6 (g)

The standard enthalpy change of reaction is given as

phpjTV1Rp.png r H 0 = phpjyELxE.pngphpXLdR9d.png H 0 f ( products ) -  phpe4m2PI.pngphp1NOs4N.png H 0 f ( reactants )

php5I6etK.pngphpSDfXlw.png H 0 f (products) = phpNmo5NR.png H 0 f C 2H6 (g) = - 84.0 kJ / mol

phpaR7SqB.png H 0f (reactants) = phpxFwLMO.png H 0 f  C2H4 (g) + php2qKz93.png H 0 f H 2(g)

phpaR7SqB.png H 0f (reactants) = phpxFwLMO.png H 0 f  C2H4 (g) + 0 kJ / mol = phpxFwLMO.png H 0 f  C2H4 (g)

From above calculated values , we can write

- 138.9 kJ / mol = ( -84.0 kJ / mol ) - phpxFwLMO.png H 0 f  C2H4 (g)

phpxFwLMO.png H 0 f  C2H4 (g) = ( -84.0 kJ / mol ) + 138.9 kJ /mol

phpxFwLMO.png H 0 f  C2H4 (g) = 54.9 kJ /mol  

ANSWER : Standard enthalpy of formation of C2H4 (g) = 54.9 kJ / mol